a lift whose cage is 3m high is moving up with an acceleration of 2m/s.a piece of stone is dropped from the top of the cage of the lift when its velocity is 8m/s.if g=10m/s square ,then the stone will reach the floor of the lift after
1
Expert's answer
2012-10-24T10:54:10-0400
h = V0 * t + (a + g) * t^2 / 2
3 = 8 * t + (2 + 10) * t^2 / 2
6 * t^2 + 8
* t - 3 = 0
t = 0.3052 s